1. (15 pts). A new warehouse is under construction that will rent out its space and material moving capabilities on a modular basis. The land for the warehouse will cost S1.5 Million, and the building will cost $4.5 Million. The building is a MACRS 39-year asset. If the warehouse will be put into service in April of 2020, and sold in November of 2028, a) compute the annual depreciation amounts and insert them into the table below for this property. b) If the Nov. 2028 sale price for land and building is $7.5 Million, compute the gains on this sale. YearDepreciation 2020 2021 2022 2023 2024 2025 2026 2027 2028
Answer #1

As per the law, Land is not depreciated because land is assumed to have an unlimited useful life. Other assets such as buildings, furnishings, equipment, etc. have limited useful lives. Therefore, they are depreciated. When an entity purchases warehouse that has land and building on it, the cost must be allocated between the land and the building; the result will be depreciation of the building, but not the land.

Here, in the question we have seperate cost for both the assets. So, building will be depreciated and not the land.

Now, MACRS depreciation schedule for a 39 year old nonresidential property is:-

Since the asset is put in the service in 4th month, rate for year 1 is 1.819%. For year 2 to 39, it is 2.564%. (You can get the table online or on IRS website)

So, Depreciation will be as follows:-

Year Depreciation

2020 45,00,000 * 1.819% = 81,855

2021 (45,000,00 - 81,855) * 2.564% = 113,281

2022 43,04,864 * 2.564% = 110,376

2023 41,94,488 * 2.564% = 107,546

2024 40,86,902 * 2.564% = 104,788

2025 39,82,114 * 2.564% = 102,101

2026 38,80,013 * 2.564% = 99,483

2027 37,80,530 * 2.564% = 96,933

2028 36,83,597 * 2.564% = 94,447

b) If the same is sold at $ 7.5 million, the gaun will be as follows,

= (Sales Value of assets - Book value of assets)

= (75,00,000 - (15,00,000 + 35,89,150))

= $ 24,10,850